Logiweb bibliography of helloworld

Up Help

[0] helloworld

[1] ijcar base

[2] ijcar

The pyk compiler, version 0.grue.20060417 by Klaus Grue GRD-2006-05-02.UTC:18:03:29.383777 = MJD-53857.TAI:18:04:02.383777 = LGT-4653309842383777e-6